If 10-7 J of heat is added to a system w is so large that its temperature () /" does not change, by what factor will the number of accessible states of the system increase if the temperature is 30 K? 


1
Expert's answer
2021-05-26T10:31:13-0400

Solution.

"dQ=7J;"

"T=30K;"

"dS=\dfrac{dQ}{T};"

"dS=\dfrac{7J}{30K}=0.23J/K;"

If "dQ=10J" then "dS=\dfrac{10J}{30K}=0.33J/K;"


Answer: "dQ=0.23 - 0.33J\/K."